# Anytime Replanning of Robot Coverage Paths for Partially Unknown Environments

### v1

@@ V The RANK-REPLAN Subroutine
In this section, we extend the mixed-integer linear program (MILP) from the OARP approach in our previous work [6] to formulate the RANK-REPLAN subroutine that minimizes the number of replanned ranks while constraining the number of new ranks within a budget $\bar{m}$. We first provide a brief review of the original MILP with some additional variables to allow for extension. Then, we propose two extensions to this MILP: the first extension solves the rank replanning step optimally, and the second extension solves it efficiently with fewer MILP variables (no optimality guarantees). The user may choose either MILP based on their replanning preferences (optimality vs efficiency).
V-A The base MILP
Consider an integral orthogonal polygon (IOP) containing $n$ grid cells representing the input environment, where each cell $c_{i}$ for $1\leq i\leq n$ is of size $l\times l$ ($l$ is the tool width). The goal of the MILP is to assign a coverage orientation (horizontal or vertical) for each cell $c_{i}$ such that the number of ranks covering the cells is minimized. We obtain ranks by merging similarly-oriented neighboring grid cells. To this end, we introduce variables to represent cell orientations and count the resulting ranks after all merges are complete.
Let $x_{H}^{i}$ and $x_{V}^{i}$ be binary variables that denote whether cell $c_{i}$ is oriented horizontally or vertically (cannot be both). Fig. 4 shows an example IOP with oriented grid cells and their corresponding ranks. We also add variables to identify the rank endpoints. Let $y_{L}^{i}$ and $y_{R}^{i}$ be binary variables that indicate whether the cell, $c_{i}$, is the left endpoint and/or the right endpoint of a horizontal rank. Similarly, let $y_{T}^{i}$ and $y_{B}^{i}$ be binary variables that indicate whether $c_{i}$ is a top and/or bottom endpoint of a vertical rank. This notation is different from the original OARP approach, which only has indicator variables for the left and top endpoints. Fig. 4 also illustrates all endpoint variables and rank endpoints for the oriented grid cells.
@@ V-A The base MILP
To formulate the MILP constraints, we must encode the relationship between the orientations ($x$ variables) and endpoints ($y$ variables). Following from our previous work [6], we use directed graphs in the IOP to detect whether a neighboring cell can be merged. Let $G_{L}$ and $G_{R}$ be directed graphs composed of all horizontal path flows for each grid cell from its left and right neighbors respectively. For cells at the IOP border, we add artificial nodes called border identifiers to simulate path flow into those cells. Similarly, let $G_{T}$ and $G_{B}$ be the graphs composed of all vertical path flows for each grid cell from its top and bottom neighbors respectively. From these graphs, let $A_{L},A_{R},A_{T},$ and $A_{B}$ be matrices representing the corresponding graphs, where each row of a matrix signifies a directed edge in the graph, with a $-1$ for a source grid cell (outgoing edge), a $+1$ for a sink grid cell (incoming edge), and 00s otherwise. These matrices resemble the node-arc incidence (NAI) matrices for the graphs $G_{L},G_{R},G_{T},$ and $G_{B}$.
We now pose the rank minimization MILP, namely MILP-0, that we will extend in the rest of this section:
MILP 0:
@@
In this MILP, equations (9) and (11) ensure that, for an optimal solution of MILP-0, vectors $\boldsymbol{y_{H}}$ and $\boldsymbol{y_{V}}$ correspond to the endpoints of the optimal number of ranks covering the IOP. Equation (12) ensures that each cell is only assigned one orientation (can be either horizontal or vertical). For an in-depth discussion about the formulation of these constraints, we refer the reader to [6].
@@ V-B Constraining new ranks
We now look to constrain the number of new ranks $m_{\text{new}}$ to be within the rank budget $\bar{m}$. To compute $m_{\text{new}}$, we identify the new ranks in the replanned solution by comparing them with the ranks from the initial path. This comparison requires an encoding of ranks using the MILP variables. To this end, we propose an encoding based on perfect matching [31] that matches the endpoints for each rank in the MILP solution. The new ranks are therefore the difference between the old and new endpoint matchings.
@@
Now, let $\bar{z}_{ij}$ be an input to the MILP such that $\bar{z}_{ij}$ is $1$ if the initial path contains a rank with cells $c_{i}$ and $c_{j}$ as endpoints and 00 otherwise. From Lemma 1, if the MILP solution has a rank with $c_{i}$ and $c_{j}$ as endpoints ($z_{ij}=1$), then it is a new rank if $\bar{z}_{ij}=0$. Using this relation for all pairs of cells $c_{i}$ and $c_{j}$ with matching edges, we compute $m_{\text{new}}$ as follows:
@@
We now propose our extended MILP, namely MILP-1, that minimizes the number of ranks in the replanned solution while constraining the number of new ranks to be within a budget $\bar{m}$. We also add the term $\epsilon\text{ }m_{\text{new}}$ to the MILP’s objective function, where $\epsilon>0$ is used to compute the rank-minimizing solution with the smallest $m_{\text{new}}$. The resulting MILP is as follows:
MILP 1:
@@
We now show that we can configure MILP-1 to solve the rank replanning stage (Problem 3(i)) optimally, i.e. the solution (i) minimizes the number of ranks $m$, (ii) subject to the constraint that $m_{\text{new}}\leq\bar{m}$.
@@
If $\epsilon<1/\bar{m}$, then MILP-1 computes optimal solutions to Problem 3(i).
@@
Any solution to MILP-1 must satisfy Eq. 26 and is therefore a feasible solution to P1. What remains to be shown is whether MILP-1 computes the solution with minimum $m$, i.e., there exists an $\epsilon>0$ such MILP-1 returns an optimizer to P1. To show this, consider the simplified version of MILP-1:
@@
The linear relaxation of MILP-1 is obtained by replacing the integrality constraints for the decision variables with non-negativity constraints. We observed that solving the relaxation consistently gives integral solutions despite the constraints not being totally unimodular (TUM). For our experiments, we simply solve the linear relaxation as LPs can be solved in polynomial time [32].
V-C On reducing MILP variables
In the previous section, we formulated MILP-1 which solves the rank replanning step optimally. However, one drawback of MILP-1 is that it adds a large number of decision variables ($\mathcal{O}(n^{2})$) to the base MILP, which affects the runtime of the MILP. In this section, we propose another MILP extension, namely MILP-2, that uses fewer decision variables and can be more successfully used for large environments. Instead of matching rank endpoints, we count new ranks by identifying changes to the rank endpoints from a previous solution. Fig. 7 illustrates all possible changes to ranks with the introduction of an obstacle. In each case, we observe changes to the rank endpoints, which we use to formulate a bound for $m_{\text{new}}$.
We first analyze the horizontal ranks to present our approach and extend the same approach to vertical ranks. Consider a binary vector $\boldsymbol{\bar{y}_{H}}\in\{0,1\}^{2n}$ that indicates if a cell was a horizontal (left or right) endpoint for a rank in the initial path. We take this vector as an input to the MILP. We look to compare vectors $\boldsymbol{y_{H}}$ and $\boldsymbol{\bar{y}_{H}}$ to identify rank endpoint changes. Specifically, we look to identify two types of changes: (i) an added endpoint and (ii) an extended endpoint.
@@ V-C On reducing MILP variables
However, in the case of a merged rank (Fig. 7), the new rank may not contain any added endpoints. To account for this, we also count extended endpoints that measure cells that were endpoints in $\boldsymbol{\bar{y}_{H}}$ and are not endpoints in $\boldsymbol{y_{H}}$ while still maintaining the same coverage orientation (horizontal in this case). Suppose a cell $c_{i}$ was a left endpoint in $\boldsymbol{\bar{y}_{H}}$, i.e. $\bar{y}_{L}^{i}=1$. The cell is considered an extended endpoint if $y_{L}^{i}=0$ and the cell is still covered horizontally ($x_{H}^{i}=1$). We use the following piece-wise equation for cell $c_{i}$ to capture this relationship.
@@
where the $\circ$ operator is the element-wise product of two vectors. We simply extend these concepts to the vertical case. Let $\boldsymbol{\bar{y}_{V}}$ be a binary vector indicating whether a cell was a vertical endpoint for a rank in the initial path. The total number of added endpoints $\alpha$ and extended endpoints $\beta$ is then obtained as follows:
@@
Figure 7: Illustration of all possible changes to ranks. Each row shows horizontal (orange lines) and vertical (purple lines) ranks blocked by obstacles (black boxes) and the corresponding replanned ranks.
@@
[…]
@@
We now propose our second MILP extension. We use the bound we derived in Lemma 2 to constrain $m_{\text{new}}$ in the MILP. Following Proposition 1, we add an additional term $\epsilon\text{ }(\alpha+\beta/2)$ to the objective function, where $\epsilon=1/(\bar{m}+1)$, to compute a solution with the smallest $m_{\text{new}}$ without dominating the minimization of $m$. The resulting MILP is as follows:
MILP 2:
@@
Solving MILP-2 guarantees a solution where $m_{\text{new}}\leq\bar{m}$. Moreover, MILP-2 only adds $\mathcal{O}(n)$ decision variables to the base MILP ($\boldsymbol{\beta_{H}}$ and $\boldsymbol{\beta_{V}}$). So, we expect MILP-2 to run faster than MILP-1. We observed that Remark 1 also applies for MILP-2, so we simply solve the linear relaxation to obtain integral solutions. However, since there are negative terms involving $\boldsymbol{y}_{H}$ and $\boldsymbol{y}_{V}$ (Eqs. (33), (35), we add an additional constraint to the linear relaxation to ensure that it is bounded:
@@
Unlike MILP-1, MILP-2 does not guarantee an optimal solution to the rank replanning problem, as we use a pessimistic bound on $m_{\text{new}}$ in our MILP constraint. However, we observe in our simulation results that the paths replanned with ranks from MILP-2 are close in cost to those from MILP-1.
@@ VI The TOUR-REPLAN Subroutine
In this section, we describe our approach to compute a tour of the replanned ranks to obtain our new coverage path. Similar to the approaches from [13] and [6], we formulate this as a Generalized Travelling Salesman Problem (GTSP) on an auxiliary graph representing all connections between ranks. Each set in the graph represents a rank and consists of two vertices to represent the directions in which the rank can be traversed. We call these sets as rank sets. The graph edges are the transition costs between the vertices, which depend on the robot’s dynamics and design. For this work, we assume robot dynamics similar to [6] and [17] where (i) the robot traverses a straight line with piecewise constant acceleration with a maximum velocity and (ii) the robot stops and turns in place with a constant angular velocity.
@@ VII Simulation Results
In this section, we evaluate the OARP-Replan approach in simulation using real-world maps. We provide coverage performance comparisons against two replanning methods: (i) greedy detour (GD) replan and (ii) OARP approach from [6] naively modified for online replanning. We also compare OARP-Replan with an offline coverage planner provided with complete obstacle information [7]. Finally, we demonstrate using OARP-Replan to replan coverage paths in a high-fidelity ROS simulation.
@@ VII-A Simulation setup
Dataset: We use a dataset of 7 real-world 2D maps obtained from Avidbots as our base environments, with their minimum coverage path lengths ranging from 145 $m$ to 1900 $m$. We then add randomly generated obstacles with varying sizes and orientations to create a set of unknown environments. Specifically, we create two datasets of unknown environments: (i) fixed clutter dataset where the new obstacles occupy 10% of the base map’s free area for all maps, and (ii) varying clutter dataset where starting with a candidate base environment, we generate unknown environments by varying the % area occupied by new obstacles. For the varying clutter dataset, we chose the base environment to be the one used in Fig. 1(a) and we test with 2 to 20 % area occupied. Due to the confidentiality of the environments, we are unable to share the complete dataset, but we have created anonymized environments for visualization, as seen in Figs. 1, 11, and 15. Fig. 11 shows an example base environment containing three unknown obstacles (red boxes) for which the initial coverage path (dark path) needs replanning.
Replan runtime estimator: To determine our rank replanning budget $\bar{m}$, we estimate the runtime $\hat{T}(m)$ needed to compute a tour for $m$ ranks. This runtime includes (i) the estimated time to compute transition paths $\hat{T}_{\text{trans}}(m)$, and (ii) the estimated time to solve the GTSP $\hat{T}_{\text{GTSP}}(m)$. Transition path computation is included as it is a time-intensive process, especially for large maps. Formally,
$\hat{T}(m)=\hat{T}_{\text{trans}}(m)+\hat{T}_{\text{GTSP}}(m).$
We use data collected from previous coverage planning runs to estimate each function. We estimate $\hat{T}_{\text{GTSP}}$ by fitting a cubic polynomial through the data as shown in Fig. 9. For $\hat{T}_{\text{trans}}(m)$, we use the following map-specific linear estimator. Let $t_{\text{avg}}$ be the average time to compute the transition paths for one rank in the given map. We compute $\hat{T}_{\text{trans}}(m)$ as follows:
$\hat{T}_{\text{trans}}(m)=t_{\text{avg}}m.$
We use the inverse of $\hat{T}(m)$ to compute the replanning budget, i.e. $\bar{m}=\hat{T}^{-1}(\tau)$. While we acknowledge that more sophisticated estimators may exist for this, we do not focus on that for this paper.
Implementation: We test the two variants of OARP-Replan (Algorithm 1) based on which RANK-REPLAN method is used: (i) OARP-Replan-1 which uses MILP-1, and (ii) OARP-Replan-2 which uses MILP-2. Both approaches were implemented in Python. For each MILP, we solve the corresponding linear relaxation using the free SCIP solver [33]. Then, we compute a tour of the replanned ranks by solving the GTSP formulation from Section VI, for which we use the open-source GLNS solver [34]. Table I shows the robot parameters used to compute the transition costs between ranks. The robot is also equipped with a radial sensor in simulation (radius in Table I) that detects new obstacles and runs OARP-Replan.
@@
Figure 9: Data used to estimate TOUR-REPLAN runtime, where we plot the number of ranks (x-axis) and the time taken to compute all transitions and replan GTSP (y-axis).
Metrics: The key metric is total coverage cost, which is the time taken by the robot to complete coverage. This is computed as a sum of drive and stoppage time, where stoppages occur when the robot is waiting for a new path to be computed. Table I shows the robot parameters used to compute the path drive time. To measure stoppage, we compute the minimum excess time taken by the algorithm to compute the replanned path as follows:
Minimum stop time $=\max(\{0,[\text{Time to plan transitions}]$
@@
VII-B Coverage cost comparisons
Figure 10: Comparison of the replanning approaches using the fixed clutter dataset: (a) Difference in coverage cost from an offline optimal planner (all obstacles known). (b) Minimum stop time of the robot during coverage (GD omitted as replanning time is negligible). The line shows the average and the lighter region shows the data interval.

### v2

@@ V The RANK-REPLAN Subroutine
In this section, we extend the integer linear program (ILP) from the OARP approach in our previous work [6] to formulate the RANK-REPLAN subroutine that minimizes the number of replanned ranks while constraining the number of new ranks within a budget $\bar{m}$. We first provide a brief review of the original ILP with some additional variables to allow for extension. Then, we propose two extensions to this ILP: the first extension solves the rank replanning step optimally, and the second extension solves it efficiently with fewer ILP variables (no optimality guarantees). The user may choose either ILP based on their replanning preferences (optimality vs efficiency).
V-A The base ILP
Consider an integral orthogonal polygon (IOP) containing $n$ grid cells representing the input environment, where each cell $c_{i}$ for $1\leq i\leq n$ is of size $l\times l$ ($l$ is the tool width). The goal of the ILP is to assign a coverage orientation (horizontal or vertical) for each cell $c_{i}$ such that the number of ranks covering the cells is minimized. We obtain ranks by merging similarly-oriented neighboring grid cells. To this end, we introduce variables to represent cell orientations and count the resulting ranks after all merges are complete.
@@ V-A The base ILP
Let $x_{H}^{i}$ and $x_{V}^{i}$ be binary variables that denote whether cell $c_{i}$ is oriented horizontally or vertically (cannot be both). Fig. 4 shows an example IOP with oriented grid cells and their corresponding ranks. We also add variables to identify the rank endpoints. Let $y_{L}^{i}$ and $y_{R}^{i}$ be binary variables that indicate whether the cell, $c_{i}$, is the left endpoint and/or the right endpoint of a horizontal rank. Similarly, let $y_{T}^{i}$ and $y_{B}^{i}$ be binary variables that indicate whether $c_{i}$ is a top and/or bottom endpoint of a vertical rank. This notation is different from the original OARP approach, which only has indicator variables for the left and top endpoints. Fig. 4 also illustrates the endpoint variables and rank endpoints for the oriented grid cells.
@@
To formulate the ILP constraints, we must encode the relationship between the orientations ($x$ variables) and endpoints ($y$ variables). Following from our previous work [6], we use directed graphs in the IOP to detect whether a neighboring cell can be merged. Let $G_{L}$ and $G_{R}$ be directed graphs composed of all horizontal path flows for each grid cell from its left and right neighbors respectively. For cells at the IOP border, we add artificial nodes called border identifiers to simulate path flow into those cells. Similarly, let $G_{T}$ and $G_{B}$ be the graphs composed of all vertical path flows for each grid cell from its top and bottom neighbors respectively. Fig. 5 illustrates these graphs for an example IOP. From these graphs, let $A_{L},A_{R},A_{T},$ and $A_{B}$ be matrices representing the corresponding graphs, where each row of a matrix signifies a directed edge in the graph, with a $-1$ for a source grid cell (outgoing edge), a $+1$ for a sink grid cell (incoming edge), and 00s otherwise. These matrices resemble the node-arc incidence (NAI) matrices for the graphs $G_{L},G_{R},G_{T},$ and $G_{B}$.
We now pose the rank minimization ILP, namely ILP-0, that we will extend in the rest of this section:
ILP-0:
@@
In this ILP, equations (8) and (9) ensure that, for an optimal solution of ILP-0, vectors $\boldsymbol{y_{H}}$ and $\boldsymbol{y_{V}}$ correspond to the endpoints of the optimal number of ranks covering the IOP. This follows from the minimization of $m$, which in-turn minimizes $\boldsymbol{y_{H}}$ and $\boldsymbol{y_{V}}$ (Eq. (6)), thereby tightening equations (8) and (9). Equation (10) ensures that each cell is only assigned one orientation (can be either horizontal or vertical). For an in-depth discussion about the formulation of these constraints, we refer the reader to [6].
Figure 5: The directed graphs $G_{L},G_{R},G_{T},$ and $G_{B}$ for the example IOP.
@@ V-B Constraining new ranks
We now look to constrain the number of new ranks $m_{\text{new}}$ to be within the rank budget $\bar{m}$. To compute $m_{\text{new}}$, we identify the new ranks in the replanned solution by comparing them with the ranks from the initial path. This comparison requires an encoding of ranks using the ILP variables. To this end, we propose an encoding based on perfect matching [35] that matches the endpoints for each rank in the ILP solution. The new ranks are therefore the difference between the old and new endpoint matchings.
@@
Now, let $\bar{z}_{ij}$ be an input to the ILP such that $\bar{z}_{ij}$ is $1$ if the initial path contains a rank with cells $c_{i}$ and $c_{j}$ as endpoints and 00 otherwise. From Lemma 1, if the ILP solution has a rank with $c_{i}$ and $c_{j}$ as endpoints ($z_{ij}=1$), then it is a new rank if $\bar{z}_{ij}=0$. Using this relation for all pairs of cells $c_{i}$ and $c_{j}$ with matching edges, we compute $m_{\text{new}}$ as follows:
@@
We now propose our extended ILP, namely ILP-1, that minimizes the number of ranks in the replanned solution while constraining the number of new ranks to be within a budget $\bar{m}$. We also add the term $\epsilon\text{ }m_{\text{new}}$ to the ILP’s objective function, where $\epsilon>0$ is used to compute the rank-minimizing solution with the smallest $m_{\text{new}}$. Specifically, in the case of a large budget $\bar{m}$, the additional term helps prevent rank changes that do not improve the number of ranks in the solution. The resulting ILP is as follows:
ILP-1:
@@
We now show that we can configure ILP-1 to solve the rank replanning stage (Problem 3(i)) optimally, i.e. the solution (i) minimizes the number of ranks $m$, (ii) subject to the constraint that $m_{\text{new}}\leq\bar{m}$.
@@
If $\epsilon<1/\bar{m}$, then ILP-1 computes optimal solutions to Problem 3(i).
@@
Any solution to ILP-1 must satisfy Eq. 24 and is therefore a feasible solution to P1. What remains to be shown is whether ILP-1 computes the solution with minimum $m$, i.e., there exists an $\epsilon>0$ such ILP-1 returns an optimizer to P1. To show this, consider the simplified version of ILP-1:
@@
The linear relaxation of ILP-1 is obtained by replacing the integrality constraints for the decision variables with real-valued constraints as follows:
$0\leq z_{ij}\leq 1\quad\forall 1\leq i,j\leq n,$
$\boldsymbol{0}\leq\boldsymbol{y_{H}},\boldsymbol{y_{V}}\leq\boldsymbol{1},$
$\boldsymbol{x_{H}},\boldsymbol{x_{V}}\geq\boldsymbol{0}.$
We observed that solving this relaxation consistently gives integral solutions despite the constraints not being totally unimodular (TUM). Our reasoning for this is that (i) the constraints of ILP-0 is TUM (following the proof from [6]), and (ii) perfect matching on a bipartite graph (e.g. graph partitioned as left endpoints ($y_{L}^{i}$ variables) and right endpoints ($y_{R}^{i}$ variables)) can be solved using a linear program (LP) [35]. For our experiments, we simply solve the linear relaxation as LPs can be solved in polynomial time [36].
V-C On reducing ILP variables
In the previous section, we formulated ILP-1 which solves the rank replanning step optimally. However, one drawback of ILP-1 is that it adds a large number of decision variables ($\mathcal{O}(n^{2})$) to the base ILP, which affects the runtime of the ILP. In this section, we propose another ILP extension, namely ILP-2, that uses fewer decision variables and can be more successfully used for large environments. Instead of matching rank endpoints to count the number of new ranks $m_{\text{new}}$, we formulate a bound for $m_{\text{new}}$ using the number of rank endpoints that have changed. Specifically, we look to identify and count two types of changes: (i) added endpoints and (ii) extended endpoints (i.e. endpoints removed when ranks from the initial solution are merged). Fig. 8 illustrates cases where the ranks are changed with the introduction of obstacles and the corresponding added and extended rank endpoints.
We first analyze the horizontal ranks to present our approach and extend the same approach to vertical ranks. Consider a binary vector $\boldsymbol{\bar{y}_{H}}\in\{0,1\}^{2n}$ that indicates if a cell was a horizontal (left or right) endpoint for a rank in the initial path. We take this vector as an input to the ILP. We look to compare vectors $\boldsymbol{y_{H}}$ and $\boldsymbol{\bar{y}_{H}}$ to identify added and extended endpoints.
@@ V-C On reducing ILP variables
However, in the case of a merged rank (Fig. 8), the new rank may not contain any added endpoints. To account for this, we also count extended endpoints that count cells that were endpoints in $\boldsymbol{\bar{y}_{H}}$ and are not endpoints in $\boldsymbol{y_{H}}$ while still maintaining the same coverage orientation (horizontal in this case). Suppose a cell $c_{i}$ was a left endpoint in $\boldsymbol{\bar{y}_{H}}$, i.e. $\bar{y}_{L}^{i}=1$. The cell is considered an extended endpoint if $y_{L}^{i}=0$ and the cell is still covered horizontally ($x_{H}^{i}=1$). We use the following piece-wise equation for cell $c_{i}$ to capture this relationship.
@@
where the $\circ$ operator is the element-wise product of two vectors. The endpoint changes that remain to be counted are the cells that were endpoints in $\boldsymbol{\bar{y}_{H}}$ and are not endpoints in $\boldsymbol{y_{H}}$ while changing coverage orientations (to vertical in this case). However, such a cell will now be a part of a vertical rank that would either have an added or extended vertical rank endpoint. Therefore, we do not count these endpoints as it tracks rank changes that are already accounted for.
We extend these concepts to the vertical case to count all endpoint changes. Let $\boldsymbol{\bar{y}_{V}}$ be a binary vector indicating whether a cell was a vertical endpoint for a rank in the initial path. The total number of added endpoints $\alpha$ and extended endpoints $\beta$ is then obtained as follows:
@@
Figure 8: Illustration of rank endpoint changes as a result of rank replanning. Each row shows horizontal (orange lines) and vertical (purple lines) ranks blocked by obstacles (black boxes) and the corresponding replanned ranks and endpoint changes.
@@
Consider $\alpha$ and $\beta$ to be the number of added and extended endpoints respectively for a set of replanned ranks. Then the number of new ranks $m_{\text{new}}$ in the set is bounded as follows:
$m_{\text{new}}\leq\alpha+\beta/2.$
@@
We now propose our second ILP extension. We use the bound we derived in Lemma 2 to constrain $m_{\text{new}}$ in the ILP. Following Proposition 1, we add an additional term $\epsilon\text{ }(\alpha+\beta/2)$ to the objective function, where $\epsilon=1/(\bar{m}+1)$, to compute a solution with the smallest $m_{\text{new}}$ without dominating the minimization of $m$. The resulting ILP is as follows:
ILP-2:
@@
Solving ILP-2 guarantees a solution where $m_{\text{new}}\leq\bar{m}$. Moreover, ILP-2 only adds $\mathcal{O}(n)$ decision variables to the base ILP ($\boldsymbol{\beta_{H}}$ and $\boldsymbol{\beta_{V}}$). So, we expect ILP-2 to run faster than ILP-1. We observed that Remark 1 also applies for ILP-2, so we simply solve the linear relaxation to obtain integral solutions. Since there are negative terms involving $\boldsymbol{y}_{H}$ and $\boldsymbol{y}_{V}$ (Eqs. (30), (31), the following constraint in the linear relaxation ensures that the optimization problem is bounded:
@@
Unlike ILP-1, ILP-2 does not guarantee an optimal solution to the rank replanning problem, as we use a pessimistic bound on $m_{\text{new}}$ in our ILP constraint. However, we observe in our simulation results that the paths replanned with ranks from ILP-2 are close in cost to those from ILP-1.
@@ VI The TOUR-REPLAN Subroutine
In this section, we describe our approach to compute a tour of the replanned ranks to obtain our new coverage path. Similar to the approaches from [14] and [6], we formulate this as a Generalized Travelling Salesman Problem (GTSP) on a fully-connected auxiliary graph representing all connections between ranks. Each set in the graph represents a rank and consists of two vertices to represent the directions in which the rank can be traversed. We call these sets as rank sets. The edge costs are the transition times between two ranks, i.e. the time for the robot to travel from one rank endpoint to another. Since the transition times depend on the robot’s design and dynamics, we compute the edge costs by assuming dynamics similar to [6] and [18], where (i) the robot traverses a straight line with piecewise constant acceleration with a maximum velocity and (ii) the robot stops and turns in place with a constant angular velocity. Solving GTSP on this graph obtains a tour that covers each rank while minimizing the transition times between the ranks. In Section VII-A, we provide more details on the robot dynamics parameters used to compute transition times in our simulations.
@@
In this paper, we created a fully-connected auxiliary graph. There are alternative approaches where the auxiliary graph need not be fully-connected, such as those using lazy edge evaluations to compute a tour [37]. However, the above challenges must still be addressed for such a formulation using the approach described for the fully-connected case.
@@ VII Simulation Results
In this section, we evaluate the OARP-Replan approach in simulation using real-world maps. We compare the performance of OARP-Replan in online coverage replanning scenarios against greedy detour (GD) replan and other planning approaches [6, 18, 32]. We also compare OARP-Replan with offline coverage planning approaches provided with complete obstacle information [7, 8], where OARP-Replan is given no prior obstacle information. Finally, we demonstrate using OARP-Replan to replan coverage paths in a high-fidelity ROS simulation.
@@ VII-A Simulation setup
Dataset: We use a dataset of 8 real-world 2D maps obtained from Avidbots as our base environments, with their minimum coverage path lengths ranging from 145 $m$ to 1900 $m$. We then add randomly generated obstacles with varying sizes and orientations to create a set of unknown environments. Specifically, we create two datasets of unknown environments: (i) fixed clutter dataset where the new obstacles occupy 10% of the base map’s free area for all maps, and (ii) varying clutter dataset where starting with a candidate base environment, we generate unknown environments by varying the % area occupied by new obstacles. For the varying clutter dataset, we chose the base environment to be the one used in Fig. 1(a) and we test with 2 to 20 % area occupied. Due to the confidentiality of the environments, we are unable to share the complete dataset, but we have created anonymized environments for visualization, as seen in Figs. 1, 12, and 16. Fig. 12 shows an example base environment containing three unknown obstacles (red boxes) for which the initial coverage path (dark path) needs replanning.
Implementation: We test the two variants of OARP-Replan (Algorithm 1) based on which RANK-REPLAN method is used: (i) OARP-Replan-1 which uses ILP-1, and (ii) OARP-Replan-2 which uses ILP-2. Both approaches were implemented in Python. For each ILP, we solve the corresponding linear relaxation using the free SCIP solver [38]. Then, we compute a tour of the replanned ranks by solving the GTSP formulation from Section VI, for which we use the open-source GLNS solver [39]. Table I shows the robot parameters used to compute the transition time between ranks. The robot is also equipped with a radial sensor in simulation (radius in Table I) that detects new obstacles and runs OARP-Replan.
Metrics: Our main metric of interest is the time taken by the robot to complete coverage, i.e. total coverage time. This is computed as a sum of drive time and stoppage time, where stoppages occur when the robot is waiting for a new path to be computed. Table I shows the robot parameters used to compute the path drive time using the dynamics model from Section VI. To measure stoppage, we compute the minimum excess time taken by the algorithm to compute the replanned path as follows:
Minimum stoppage time $=\max(\{0,[\text{Time to plan transitions}]$
@@
Figure 10: Data used to estimate TOUR-REPLAN runtime, where we plot the number of ranks (x-axis) and the time taken to compute all transitions and replan GTSP (y-axis). The red line shows the cubic polynomial fit to the data.
Replan runtime estimator: In each simulation run, the robot conducts many replans of the coverage path depending on the obstacles it observes. For each replan, we compute a time $\tau$ to reach the next obstacle encounter along the current path using the robot model from Section VI. Using $\tau$, we then determine a budget $\bar{m}$ for rank replanning using an estimator function $\hat{T}(m)$ of the runtime needed to compute a tour for $m$ ranks. The inverse of $\hat{T}(m)$ is then used to compute the replanning budget, i.e. $\bar{m}=\hat{T}^{-1}(\tau)$.
The runtime estimate $\hat{T}(m)$ includes (i) the estimated time to compute transition paths $\hat{T}_{\text{trans}}(m)$, and (ii) the estimated time to solve the GTSP $\hat{T}_{\text{GTSP}}(m)$. Transition path computation is included as it is a time-intensive process, especially for large maps. Formally,
$\hat{T}(m)=\hat{T}_{\text{trans}}(m)+\hat{T}_{\text{GTSP}}(m).$
We use data collected from previous coverage planning runs to estimate each function. Our solver of choice, GLNS, typically computes a tour in $\mathcal{O}(m^{2})$ time for $m$ ranks [39]. Therefore, we estimate $\hat{T}_{\text{GTSP}}$ by fitting a cubic polynomial through the data as shown in Fig. 10, where we plot the time to compute the tour for different rank counts ($m$ values). For $\hat{T}_{\text{trans}}(m)$, we use the following map-specific linear estimator. Let $t_{\text{avg}}$ be the average time to compute the transition paths for one rank in the given map. We compute $\hat{T}_{\text{trans}}(m)$ as follows:
$\hat{T}_{\text{trans}}(m)=t_{\text{avg}}m.$
Under-estimating the replan runtime will result in robot stoppage, as the robot must wait for the new path to be computed. Using the above estimator with OARP-Replan in our simulations, we observed little to no such robot stoppage. Therefore, we use this simple estimator for the rest of the paper as it works well for OARP-Replan. Regardless, this estimator may be inaccurate in cases where the time to compute transition path differs significantly from $t_{\text{avg}}$ (e.g. time to compute transitions increase when a large number of obstacles are added to an environment). While other sophisticated estimators can be used with OARP-Replan to handle this, we do not focus on this for this paper.
VII-B Comparisons with online approaches
In this section, we compare the total coverage time (drive time $+$ stop time) of our two proposed OARP-Replan approaches against online coverage planning/replanning approaches. Our baseline approaches are as follows: (i) the OARP approach as proposed in [6], (ii) the heuristic (HR) approach from [18], (iii) the online search-based (SB) coverage planning approach from [32], and (iv) GD replan, i.e. the practical case of following a turn-minimizing coverage path with a local planner. The first two approaches minimize turns in the path but they are offline planners. So we modify them to be usable online for replanning as follows:
  1. (i)
    while the robot is following the path and detects new obstacle encounters, the planning approach is used to replan the ranks for the remaining coverage path, and
  2. (ii)
    the tour of the resulting ranks is computed by solving the GTSP as formulated in Section VI, where unchanged parts of the coverage path are preserved.
For OARP, we replan coverage ranks by solving the linear relaxation of ILP-0 from Section V (identical to the LP in [6]). We refer to this online variant of OARP as OARP-Original. For the HR approach, we solve Algorithm 1 from [18] to replan ranks. The latter two approaches (SB and GD) are designed for online coverage planning with fast computation of coverage paths. As a result, these approaches have negligible stoppage time. The online SB approach plans the entire coverage path online and does not require an initial path. However, the SB approach minimizes path length rather than total coverage time with turns.
Figure 11: Comparison of the online replanning approaches using the fixed clutter dataset: (a) Difference between total coverage time and base coverage time. (b) Minimum stop time of the robot during coverage (GD and SB omitted as replanning time is negligible). The line shows the average and the lighter region shows the data interval.